what is the factored form of x2y+3yx-4y?
Answer by stanbon(75887)

factored form of x2y+3yx-4y
-----
x^2y + 3xh - 4y
----
Common Factor: y
----
= y(x^2 + 3x - 4)
----
= y(x^2 +4x-x -4)
----
= y(x(x+4)-(x+4))
----
= y(x+4)(x-1)
------------------
Cheers,
Stan H.